# Performance of Superconducting Resonators Suspended on SiN Membranes

### v1

@@ Appendix A ADR Setup
Most characterization of devices was performed in an HPD Rainier 103 adiabatic demagnetization refrigerator (ADR) at a base temperature of 100 mK. The samples were measured in reflection using a central bus to interact with all resonators, minimizing the needed number of lines. A VNA was used for all measurements, and the ADR temperature could be reliably modified to perform temperature sweeps. Further details on the microwave setup are presented in Fig. 5, similar to past work where resonator characterization was performed [38].
Appendix B Fabrication Procedure
@@
While the main fabrication steps were outlined in Fig. 2(a), we provide more extensive details here. To begin, a 6 inch, 675 m thick high-resistivity ( ohm-cm) Si wafer was cleaned using a piranha etch and a buffered-oxide etch (BOE). A 110 nm layer of Si-rich was then grown on both sides of the wafer [42]. Following this, 125 nm of Nb was sputtered on one side. A conventional optical lithography procedure was performed to write and develop the ground plane features. The Nb was then etched to yield these completed features, which were protected before moving on to write the etch windows on the opposite side of the wafer. Once developed, these windows were used to selectively etch through the SiN layer and Si wafer beneath the resonators using consecutive anisotropic, deep reactive ion etching (DRIE) following the Bosch process, with a target total etching depth of 600 m. Due to the fragility of the membranes but necessity of galvanically stitching together the ground plane, airbridges were written using a single-step grayscale process. After ion milling to ensure galvanic contact, Al was evaporated to yield the airbridges.
@@
Liftoff following this evaporation was delayed due to the airbridge fragility; instead, the wafer topside was protected with a dedicated photoresist before etching the remaining Si with KOH, which self-terminates at the SiN layer. This overall two-part etch series of a dry then wet etch was challenging to optimize. Use of only KOH instead of starting with a dry etch was unsuccessful due to the etch spreading too wide while etching through the Si wafer, rendering it challenging to achieve target membrane dimensions. We found that beginning with the dry, anisotropic etch was necessary. Another component to optimize revolved around how KOH also etches Nb. Consequently, any exposed feature or side of the Nb layer would lead to a wafer-scale failure during the wet etch. Thus, in addition to the protective layer, a clamping tool was used to further isolate the wafer topside from KOH while etching.
@@
With the membranes now formed but the wafer still whole, individual dies were cleaved by hand, following cleaving lines formed earlier. Using a dicing saw would be too vigorous and destroy the membranes. Once an individual die was obtained, liftoff was performed using room temperature acetone overnight, for around 12 hours. This differs from other liftoff techniques using hot (67∘C) acetone or hot (80∘C) or even room temperature Microposit Remover 1165, which were all too aggressive for the membranes [57]. Finally, the dies were transferred to IPA, where the solvent was removed using a critical point dryer (CPD). Again, removal of solvent by other means such as blow drying, spin drying, or a hot plate destroyed the membranes.
Appendix C Materials Characterization
@@
Materials characterization measurements were performed to examine the quality of the devices, similar to past work [38]. In preparation for TEM imaging, cross-section lamellae were prepared by FIB machining and extracted from the resonator chips. They were prepared at the Surface Analysis Laboratory, UTAH NANOFAB using a FEI Helios Nanolab 650 dual-beam at the University of Utah. Once the samples were received back to LBNL, TEM images and EDS elemental maps, such as that in Fig. 2 (b), were acquired at 200 kV using a JEOL 2100-F field emission scanning transmission electron microscope (STEM), which was equipped with an Oxford high-solid-angle silicon drift detector (SDD) x-ray energy dispersive spectrometer. For EDS spectral imaging, a 1-nm-diameter electron-beam probe was used. For finer resolution measurements, such as that in Fig. 2 (d), STEM-EELS was performed at 300 kV using the TEAM I double-aberration-corrected STEM equipped with a high-resolution Continuum Gatan imaging filter (GIF) spectrometer and a 4K 4K Gatan K3 direct electron detector. EELS spectral imaging maps were acquired using an electron-beam probe of a 0.1 nm diameter with 1-eV elastic-peak energy-resolution. A measurement of this chemical mapping is displayed in Fig. 6, showing the clean interface between the Si and SiN layers.
@@
The resonator characterization is also notable in light of the Nb being grown on amorphous SiN, in contrast to the more standard case of growth on single crystalline Si. The Nb film on SiN exhibits a smaller grain size with random crystallographic orientation as compared to the larger grain sizes and preferred crystallographic orientation with apparent epitaxial Nb-grain nucleation observed in Nb on Si [38]. These distinctions raise further avenues of exploration concerning Nb growth on amorphous substrates and how the differences in grain size and orientation could affect its material properties and chemistry.
Appendix D Resonator Details
@@
Here, we provide the measured values for each resonator of the two samples, measured at low power (single photon level) and high power, as shown in Fig. 7. Each chip had four resonators of each type on it. Although all resonators were found on Sample B, instead six were observed on Sample A.
@@
Regarding the fitting routine of the resonator response versus power, in light of more noisy data at low power values, all fits were restricted to the interval above . Points were defined as outliers, likely due to noise at low powers or abnormal effects at high powers as also seen in other work [58, 59], and removed from the fitting procedure starting at if: (1) any low power point had a that decreased by more than relative to its neighboring next point, and (2) any high power point had a that increased by more than relative to its neighboring previous point.
Appendix E Airbridge Effect Modelling
@@
The observed small relative change of frequency () with temperature () suggests that -dependent effects can be considered as perturbations to the resonant mode. In this cavity-perturbation regime [60], the evolution of the complex resonant frequency follows
@@
We use the following expression for the TLS contribution [31, 43]:
@@
For the airbridge contribution, as the penetration depth in Al is much smaller than the coherence length [61] and the airbridge thickness (615 nm as measured by TEM at the pads in contact with Nb, see Fig. 8), we calculate using
and according to the Mattis–Bardeen theory [48],
Here is the conductivity in the normal state, is the vacuum permeability, is the Fermi function at , and the superconducting gap is approximated as with representing the zero-temperature gap size. Eqs. (6) and (7) are valid for , which always holds in the temperature and frequency range of our measurements.
As shown in Fig. 9, we fit the real and imaginary parts of simultaneously using Eqs. (2)-(7). The independent fitting parameters are , , , , and . The integrals in Eqs. (6) and (7) are evaluated numerically using code modified from Ref. [62]. The fit well reproduces the overall temperature dependence, with mild overestimation of hardly visible on the plotted frequency scale. The surface impedance term is found to dominate the response, and the extracted and are in reasonable agreement with literature values for Al [61, 47]. These results indicate that the temperature-dependent effects can mainly be attributed to the Al airbridges.
Appendix F Thermalization Rate Comparison Analysis
Here, we present further information on the comparisons between resonator thermalization rates. For each region of interest, shaded in different colors in Fig. 10, the two resonator frequencies versus recorded point (earlier converted to temperature versus time in Fig. 4) were examined more closely to determine any significant differences. R1 and R4 are on SiN membranes, while R3, R5, and R6 are on-substrate.
In region I (III), resonator thermalization differences would result in one resonator’s frequency changing sooner than the other. To identify the point when the resonator’s frequency started to change, with noise in mind, a moving average using a window of points was run across the data. The difference was taken between each window’s mean and the prior window’s, and the resonator’s initial point of change was identified as the point when all further window differences were negative (positive), meaning that the moving averages continually became lower (higher). In region II (IV), the resonator frequency versus point was fit to an asymmetric parabola, and the minimum (maximum) points were compared. The difference in these corresponding points for each pair of resonators is presented in Table 1.
@@ Appendix F Thermalization Rate Comparison Analysis
Appendix G Frequency Stability
To explore any differences in the temporal frequency stability of the resonators, a single point from each was monitored over approximately 90 minutes using a VNA at . A single point was chosen instead of a full scan over the resonator to achieve a faster repetition rate. Using initial scans across all resonators, by fitting the resonator’s phase versus frequency in an interval close to resonance, a linear relationship can be extracted to convert between the two [63]. This linear relationship is formally expressed by the Barkhausen relationship, [64, 65].
This calibration and measurement procedure was performed for on-substrate and membrane resonators on the same chip, as shown in Fig. 11. In contrast to results presented earlier, these measurements were performed in a dilution refrigerator at a temperature of 27 mK. To compare the resonator stability, we calculated the standard deviations () of the extracted frequency values over the duration of the scans [66], with results of each resonator presented in Table 2. Although certain resonators were more stable than others, no distinction was observed between membrane versus on-substrate resonators. Therefore, these results suggest that fabricating resonators on thin SiN membranes does not adversely affect their temporal frequency stability.
@@ Appendix G Frequency Stability
Appendix H HFSS Simulations
@@
To address this concern and verify the compatibility between qubits and these membranes, we conducted simulations using Ansys HFSS, a commonly used software for simulating superconducting circuit designs. We found that adding “fingers" to the qubit capacitor paddles yielded the desired capacitance ( fF); furthermore, moving the resonator coupling section closer to the qubit paddle and within the gap between it and the ground plane also yielded the desired qubit-resonator coupling ( MHz), both values typical of other transmon designs [67]. Both of these simulations were performed using the Eigenmode solver of HFSS. To identify the qubit capacitance, a lumped element inductor was placed between the capacitor paddles, modelled as 2D sheets with a perfect E boundary condition. From the simulated frequency of the associated mode, the corresponding capacitance was calculated. Meanwhile, was simulated by sweeping the inductance of the qubit mode through the resonator mode and scaling the avoided crossing difference back to the frequencies of interest. With these simulations as guidance, one can begin to explore placing qubits on SiN membranes.

### v2

@@ I ADR Setup
Most characterization of devices was performed in an HPD Rainier 103 adiabatic demagnetization refrigerator (ADR) at a base temperature of 100 mK. The samples were measured in reflection using a central bus to interact with all resonators, minimizing the needed number of lines. A VNA was used for all measurements, and the ADR temperature could be reliably modified to perform temperature sweeps. Further details on the microwave setup are presented in Fig. S1, similar to past work where resonator characterization was performed Altoé et al. (2022).
II Fabrication Procedure
@@
While the main fabrication steps were outlined in the main text, we provide more extensive details here. To begin, a 6 inch, 675 m thick high-resistivity ( ohm-cm) Si wafer was cleaned using a piranha etch and a buffered-oxide etch (BOE). A 110 nm layer of Si-rich was then grown on both sides of the wafer Paik and Osborn (2010). Following this, 125 nm of Nb was sputtered on one side. A conventional optical lithography procedure was performed to write and develop the ground plane features. The Nb was then etched to yield these completed features, which were protected before moving on to write the etch windows on the opposite side of the wafer. Once developed, these windows were used to selectively etch through the SiN layer ( based) and Si wafer ( based) beneath the resonators using consecutive anisotropic, deep reactive ion etching (DRIE) following the Bosch process, with a target total etching depth of 600 m. Due to the fragility of the membranes but necessity of galvanically stitching together the ground plane, airbridges were written using a single-step grayscale process. After development and ion milling to ensure galvanic contact, Al was evaporated to yield the airbridges.
@@
Liftoff following this evaporation was delayed due to the airbridge fragility; instead, the wafer topside was protected with a dedicated photoresist before etching the remaining Si with KOH, which self-terminates at the SiN layer. This overall two-part etch series of a dry then wet etch was challenging to optimize. Use of only KOH instead of starting with a dry etch was unsuccessful due to the etch spreading too wide while etching through the Si wafer, rendering it challenging to achieve target membrane dimensions. We found that beginning with the dry, anisotropic etch was necessary. Another component to optimize revolved around how KOH also etches Nb. Consequently, any exposed feature or side of the Nb layer would lead to a wafer-scale failure during the wet etch. Thus, in addition to the protective layer, a clamping tool with an O-ring was used to only expose the etch windows to the KOH.
@@
With the membranes now formed but the wafer still whole, individual dies were cleaved by hand, following cleaving lines formed earlier. Using a dicing saw would be too vigorous and break the membranes. Once an individual die was obtained, liftoff was performed using room temperature acetone overnight, for around 12 hours. This differs from other liftoff techniques using hot (67∘C) acetone or hot (80∘C) or even room temperature Microposit Remover 1165, which were all too aggressive for the membranes Kreikebaum et al. (2020). Finally, the dies were transferred to IPA, where the solvent was removed using a critical point dryer (CPD). Again, removal of solvent by other means such as blow drying, spin drying, or a hot plate broke the membranes. After this extensive optimization, the ultimate yield of a chip with all four membranes intact was approximately , with the final CPD step having the lowest success rate. However, utilizing a compatible sample holder for these chips could improve yield by keeping the chip positions more fixed during the CPD process.
III Materials Characterization
@@
Materials characterization measurements were performed to examine the quality of the devices, similar to past work Altoé et al. (2022). In preparation for TEM imaging, cross-section lamellae were prepared by FIB machining and extracted from the resonator chips. They were prepared at the Surface Analysis Laboratory, UTAH NANOFAB using a FEI Helios Nanolab 650 dual-beam at the University of Utah. Once the samples were received back to LBNL, TEM images and EDS elemental maps were acquired at 200 kV using a JEOL 2100-F field emission scanning transmission electron microscope (STEM), which was equipped with an Oxford high-solid-angle silicon drift detector (SDD) x-ray energy dispersive spectrometer. For EDS spectral imaging, a 1-nm-diameter electron-beam probe was used. For finer resolution measurements, STEM-EELS was performed at 300 kV using the TEAM I double-aberration-corrected STEM equipped with a high-resolution Continuum Gatan imaging filter (GIF) spectrometer and a 4K 4K Gatan K3 direct electron detector. EELS spectral imaging maps were acquired using an electron-beam probe of a 0.1 nm diameter with 1-eV elastic-peak energy-resolution. A measurement of this chemical mapping is displayed in Fig. S2, showing the clean interface between the Si and SiN layers.
@@
The resonator characterization is also notable in light of the Nb being grown on amorphous SiN, in contrast to the more standard case of growth on single crystalline Si. The Nb film on SiN exhibits a smaller grain size with random crystallographic orientation as compared to the larger grain sizes and preferred crystallographic orientation with apparent epitaxial Nb-grain nucleation observed in Nb on Si Altoé et al. (2022). These distinctions raise further avenues of exploration concerning Nb growth on amorphous substrates and how the differences in grain size and orientation could affect its material properties and chemistry.
IV Resonator Details
@@
Here, we provide the measured values for each resonator of the two samples, measured at low power (single photon level) and high power, as shown in Fig. S3. Each chip had four resonators of each type on it. Although all resonators were found on Sample B, instead six were observed on Sample A.
@@
Regarding the fitting routine of the resonator response versus power, in light of more noisy data at low power values, all fits were restricted to the interval above . Points were defined as outliers, likely due to noise at low powers or abnormal effects at high powers as also seen in other work Crowley et al. (2023); Drimmer et al. (2024), and removed from the fitting procedure starting at if: (1) any low power point had a that decreased by more than relative to its neighboring next point, and (2) any high power point had a that increased by more than relative to its neighboring previous point.
V Airbridge Effect Modelling
@@
The observed small relative change of frequency () with temperature () suggests that -dependent effects can be considered as perturbations to the resonant mode. In this cavity-perturbation regime Klein et al. (1993), the evolution of the complex resonant frequency follows
@@
We use the following expression for the TLS contribution Gao (2008); Pappas et al. (2011):
@@
For the airbridge contribution, as the penetration depth in Al is much smaller than the coherence length Miller (1960) and the airbridge thickness (615 nm as measured by TEM at the pads in contact with Nb, see Fig. S4), we calculate using
and according to the Mattis–Bardeen theory Mattis and Bardeen (1958),
Here is the conductivity in the normal state, is the vacuum permeability, is the Fermi function at , and the superconducting gap is approximated as with representing the zero-temperature gap size. Eqs. (S5) and (S6) are valid for , which always holds in the temperature and frequency range of our measurements.
As shown in Fig. S5, we fit the real and imaginary parts of simultaneously using Eqs. (S1)-(S6). The independent fitting parameters are , , , , and . The integrals in Eqs. (S5) and (S6) are evaluated numerically using code modified from Ref. Aude (2010). The fit well reproduces the overall temperature dependence, with mild overestimation of hardly visible on the plotted frequency scale. The surface impedance term is found to dominate the response, and the extracted and are in reasonable agreement with literature values for Al Miller (1960); Caplan and Chanin (1965). These results indicate that the temperature-dependent effects can mainly be attributed to the Al airbridges.
Notably, the airbridge impact on the resonator and frequency offers a temperature bound beyond which Al airbridges should no longer be used, with values decreasing to half their maximum at around mK. This property is important to keep in mind as efforts continue to explore operating superconducting circuits at higher temperatures Anferov et al. (2023, 2024).
VI Thermalization Rate Comparison Analysis
Here, we present further information on the comparisons between resonator thermalization rates. For each region of interest, shaded in different colors in Fig. S6, the two resonator frequencies versus recorded point were examined more closely to determine any significant differences. R1 and R4 are on SiN membranes, while R3, R5, and R6 are on-substrate.
In region I (III), resonator thermalization differences would result in one resonator’s frequency changing sooner than the other. To identify the point when the resonator’s frequency started to change, with noise in mind, a moving average using a window of points was run across the data. The difference was taken between each window’s mean and the prior window’s, and the resonator’s initial point of change was identified as the point when all further window differences were negative (positive), meaning that the moving averages continually became lower (higher). In region II (IV), the resonator frequency versus point was fit to an asymmetric parabola, and the minimum (maximum) points were compared. The difference in these corresponding points for each pair of resonators is presented in Table S1.
@@ VI Thermalization Rate Comparison Analysis
VII Frequency Stability
To explore any differences in the temporal frequency stability of the resonators, a single point from each was monitored over approximately 90 minutes using a VNA at . A single point was chosen instead of a full scan over the resonator to achieve a faster repetition rate. Using initial scans across all resonators, by fitting the resonator’s phase versus frequency in an interval close to resonance, a linear relationship can be extracted to convert between the two Grünhaupt et al. (2018). This linear relationship is formally expressed by the Barkhausen relationship, Rubiola (2010); Niepce (2020).
This calibration and measurement procedure was performed for on-substrate and membrane resonators on the same chip, as shown in Fig. S7. In contrast to results presented earlier, these measurements were performed in a dilution refrigerator at a temperature of 27 mK. To compare resonator stability, we calculated the standard deviations () of the extracted frequency values over the duration of the scans Niepce et al. (2021), with results of each resonator presented in Table S2. Although certain resonators were more stable than others, no distinction was observed between membrane versus on-substrate resonators. Therefore, these results suggest that fabricating resonators on thin SiN membranes does not adversely affect their temporal frequency stability.
@@ VII Frequency Stability
VIII HFSS Simulations
@@
To address this concern and verify the compatibility between qubits and these membranes, we conducted simulations using Ansys HFSS, a commonly used software for simulating superconducting circuit designs. We found that adding “fingers" to the qubit capacitor paddles yielded the desired capacitance ( fF); furthermore, moving the resonator coupling section closer to the qubit paddle and within the gap between it and the ground plane also yielded the desired qubit-resonator coupling ( MHz), both values typical of other transmon designs Nguyen et al. (2023). Both of these simulations were performed using the Eigenmode solver of HFSS. To identify the qubit capacitance, a lumped element inductor was placed between the capacitor paddles, modelled as 2D sheets with a perfect E boundary condition. From the simulated frequency of the associated mode, the corresponding capacitance was calculated. Meanwhile, was simulated by sweeping the inductance of the qubit mode through the resonator mode and scaling the avoided crossing difference back to the frequencies of interest. With these simulations as guidance, one can begin to explore placing qubits on SiN membranes.